# Quantum adiabatic optimization with Rydberg arrays: localization phenomena and encoding strategies

### v1

@@ H.2 Positioning of the ancillae $B$
In the perturbative limit, the extended two-domain-wall subspace (the lowest energy subspace of $H_{0}$) is now also labeled by the occupation number of $B_{w}$: $|j_{x},j_{y};n_{A_{x}},n_{A_{y}};n_{B_{x}},n_{B_{y}}\rangle$. We strategically position the $B_{x}$ ancilla in a way that it can only be excited for $j_{x}=\zeta_{x}+2$, $j_{y}=\zeta_{y}$, $n_{A_{x}}=1$ and $n_{A_{y}}=0$. This effectively creates a copy, labeled by $|{\zeta_{x}+2},\zeta_{y};1,0;1,0\rangle$, of the corner state $|\zeta_{x},{\zeta_{y}+2};1,0;0,0\rangle$. Symmetrically, $B_{y}$ generates a copy, labeled by $|\zeta_{x},{\zeta_{y}+2};0,1;0,1\rangle$, of the other corner state $|\zeta_{x},{\zeta_{y}+2};0,1;0,0\rangle$.
@@ Appendix I Tackling exponential gaps via boundary Rabi frequency control
Figure 23: Boundary Rabi frequency control method. (a) Fidelity susceptibility [Eq. (3) with $\delta\lambda=0.005$] along the standard protocol for the configuration in Fig. 4(b) with $N=100$ atoms per vertex-wire, setting $\delta_{x}/\Delta_{0}=0.6$ and $\delta_{y}/\Delta_{0}=0$. We compare the two cases of (orange) homogeneous Rabi frequency and of (blue) applying the boundary Rabi frequency control method. The dashed gray lines indicate the transition point known for one-dimensional Rydberg chains with nearest-neighbor blockade, $\Omega/\Delta\sim 1/1.3$ [49], and the point where the fidelity is maximum given homogeneous Rabi frequency, $(\Delta/\Omega)_{\rm max}\sim 1.5$. (b) Position of the maximum of $\chi$, $\left(\Delta/\Omega\right)_{\rm max}$, along the standard protocol as a function of the wire weight $\delta_{x}/\Delta_{0}$ (for $\delta_{y}=0$), with $N=60$. We compare the two cases of homogeneous Rabi frequency ($\delta\lambda=0.005$) and of applying the boundary Rabi frequency control method ($\delta\lambda=0.0005$). (c) Minimum gap scaling as a function of $N$ for the four configurations of Fig. 4 after applying the boundary Rabi frequency control method along the standard protocol. For the configuration in Fig. 4(b), the data corresponding to the two different final ground states $|01\rangle_{l}$ and $|10\rangle_{l}$ are shown in orange and blue, respectively.
Here, we present a strategy to mitigate exponentially closing gaps, making use of local control of the Rabi frequency at the boundary of the vertex-wires. This approach is driven by numerical observations in the standard protocol.
We illustrate the method for the (b) configuration of Fig. 4 when targeting the $|10\rangle_{l}$ state. We first discuss the ground state properties along the standard protocol. As $\Delta/\Omega$ increases above 1.3, the ground state continuously transitions from the disordered to the ordered phase. Within the ordered phase, we detect a dominant peak in the fidelity susceptibility at around $(\Delta/\Omega)_{\rm max}\sim 1.5$ (where $\chi\simeq 1500$), corresponding to the exponentially closing gap dominating the adiabatic dynamics [see Fig. 23(a)].

### v2

@@ H.2 Positioning of the ancillae $B$
In the perturbative limit, the extended two-domain-wall subspace (the lowest-energy subspace of $H_{0}$) is now also labeled by the occupation number of $B_{w}$: $|j_{x},j_{y};n_{A_{x}},n_{A_{y}};n_{B_{x}},n_{B_{y}}\rangle$. We strategically position the $B_{x}$ ancilla in a way that it can only be excited for $j_{x}=\zeta_{x}+2$, $j_{y}=\zeta_{y}$, $n_{A_{x}}=1$ and $n_{A_{y}}=0$. This effectively creates a copy, labeled by $|{\zeta_{x}+2},\zeta_{y};1,0;1,0\rangle$, of the corner state $|\zeta_{x},{\zeta_{y}+2};1,0;0,0\rangle$. Symmetrically, $B_{y}$ generates a copy, labeled by $|\zeta_{x},{\zeta_{y}+2};0,1;0,1\rangle$, of the other corner state $|\zeta_{x},{\zeta_{y}+2};0,1;0,0\rangle$.
@@ Appendix I Tackling exponential gaps via boundary Rabi frequency control
Figure 23: Boundary Rabi frequency control method. (a) Fidelity susceptibility [Eq. (3) with $\delta\lambda=0.005$] along the standard protocol for the configuration in Fig. 4(b) with $N=100$ atoms per vertex wire, setting $\delta_{x}/\Delta_{0}=0.6$ and $\delta_{y}/\Delta_{0}=0$. We compare the two cases of (orange) homogeneous Rabi frequency and of (blue) applying the boundary Rabi frequency control method. The dashed gray lines indicate the transition point known for one-dimensional Rydberg chains with nearest-neighbor blockade, $\Omega/\Delta\sim 1/1.3$ [51], and the point where the fidelity is maximum given homogeneous Rabi frequency, $(\Delta/\Omega)_{\rm max}\sim 1.5$. (b) Position of the maximum of $\chi$, $\left(\Delta/\Omega\right)_{\rm max}$, along the standard protocol as a function of the wire weight $\delta_{x}/\Delta_{0}$ (for $\delta_{y}=0$), with $N=60$. We compare the two cases of homogeneous Rabi frequency ($\delta\lambda=0.005$) and of applying the boundary Rabi frequency control method ($\delta\lambda=0.0005$). (c) Minimum gap scaling as a function of $N$ for the four configurations of Fig. 4 after applying the boundary Rabi frequency control method along the standard protocol. For the configuration in Fig. 4(b), the data corresponding to the two different final ground states $|01\rangle_{l}$ and $|10\rangle_{l}$ are shown in orange and blue, respectively.
Here, we present a strategy to mitigate exponentially closing gaps, making use of local control of the Rabi frequency at the boundary of the vertex wires. This approach is driven by numerical observations in the standard protocol.
We illustrate the method for the (b) configuration of Fig. 4 when targeting the $|10\rangle_{l}$ state. We first discuss the ground-state properties along the standard protocol. As $\Delta/\Omega$ increases above 1.3, the ground-state continuously transitions from the disordered to the ordered phase. Within the ordered phase, we detect a dominant peak in the fidelity susceptibility at around $(\Delta/\Omega)_{\rm max}\sim 1.5$ (where $\chi\simeq 1500$), corresponding to the exponentially closing gap dominating the adiabatic dynamics [see Fig. 23(a)].